OpenStudy (anonymous):

Will award medal!!!! find a) the curl and b) the divergence of the vector field.

OpenStudy (anonymous):

\[F(x,y,z,)=\frac{ 1 }{ \sqrt(x^2+y^2+z^2 )} (x i+yj+zk)\]

OpenStudy (anonymous):

@iambatman

OpenStudy (perl):

we can solve this using a determinant

OpenStudy (anonymous):

can u guys help me plz http://openstudy.com/study#/updates/55366a0ee4b0634ee6726953

OpenStudy (perl):

$$ \Large {\\ {\bf F}(x,y,z,)=\frac{(x {\bf i}+y{\bf j}+z{\bf k})}{ \sqrt{x^2+y^2+z^2 }} \\~\\ \\ curl {~\bf F} =\nabla \times {~\bf F} \\= \begin{vmatrix} {\bf i} & {\bf j}& {\bf k}\\ \frac{\partial }{\partial x } &\frac{\partial }{\partial y } & \frac{\partial }{\partial z } \\\frac{x }{ \sqrt{x^2+y^2+z^2 }} & \frac{y }{ \sqrt{x^2+y^2+z^2 }} &\frac{z }{ \sqrt{x^2+y^2+z^2 }} \end{vmatrix} }$$

OpenStudy (irishboy123):

\[\vec F (x,y,z) = \frac{1}{\sqrt{\rho}}<x,y,z>; \ \ \rho = x^2 + y^2 + z^2\] CURL F is conservative if there exists scalar function V such that \[\vec F = \vec{\nabla} V\] \[\vec{\nabla} (\rho^{\frac{1}{2}}) = <\frac{1}{2} \rho^{-\frac{1}{2}}(2x),\frac{1}{2} \rho^{-\frac{1}{2}}(2y),\frac{1}{2} \rho^{-\frac{1}{2}}(2z)>\] \[\vec{F} = \vec{\nabla} (\rho^{\frac{1}{2}}) = \frac{1}{\sqrt{\rho}}<x,y,z>\] F is conservative so curl F = 0. DIV \[\nabla \bullet \vec{F} = \frac{∂}{∂x}\frac{x}{\sqrt{\rho}} + \frac{∂}{∂y}\frac{y}{\sqrt{\rho}} + \frac{∂}{∂z}\frac{z}{\sqrt{\rho}}\] \[\frac{∂}{∂x}\frac{x}{\sqrt{\rho}} = \frac{\sqrt{\rho} - x\frac{x}{\sqrt{\rho}} }{\rho} = \frac{1}{\sqrt{\rho}} - \frac{x^2}{\rho^{\frac{3}{2}}}\] \[\vec{\nabla} \bullet \vec{F} = \frac{3}{\sqrt{\rho}} - \frac{x^2 + y^2 + z^2}{\rho^\frac{3}{2}} = \frac{2}{\sqrt{x^2 + y^2 + z^2}}\]